# Birch SGD: A Tree Graph Framework for Local and Asynchronous SGD Methods

### v2

@@ 1 Introduction
Optimization is central to machine learning (ML), data science (DS), and federated learning (FL) (Konečný et al.,, 2016; Bottou et al.,, 2018; Kairouz et al.,, 2021). In these domains, stochastic optimization techniques such as stochastic gradient descent (SGD) (Robbins and Monro,, 1951) and its adaptive variants (Adam, AdamW, etc) (Kingma and Ba,, 2015; Loshchilov and Hutter,, 2019) have become the standard approach for tackling large-scale problems (Schmidt et al.,, 2021). Due to the rising computational demands of modern functions, the theoretical foundation of distributed algorithms supporting a large number of workers (e.g., CPUs, GPUs, servers) is important (Mayer and Jacobsen,, 2020; Kairouz et al.,, 2021; Douillard et al.,, 2023).
@@
In nonconvex settings, the goal is to find an $\varepsilon$–stationary point, meaning we want to find a random vector $\bar{x}$ such that $\mathbb{E}\left[\|\nabla f(\bar{x})\|^{2}\right]\leq\varepsilon$ (Nemirovskij and Yudin,, 1983; Murty and Kabadi,, 1985). The function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies the following standard assumptions:
@@
$f$ is differentiable and $L$–smooth: $\left\|\nabla f(x)-\nabla f(y)\right\|\leq L\left\|x-y\right\|$ $\forall x,y\in\mathbb{R}^{d}.$
@@
We focus on problems where workers are limited to computing stochastic gradients. Each worker has access to unbiased stochastic gradients, denoted by $\nabla f(x;\xi)$, whose variance is bounded by $\sigma^{2}$. In the context of ML, this implies that all workers can access the same data, which is practical when training large language and computer vision models. In such scenarios, privacy is not a critical concern, and devices can sample data from the Internet or shared datasets (Goodfellow et al.,, 2016).
@@ 1.1 Related work
One worker and optimal oracle complexity. With a single worker, the most standard optimization method is the Vanilla SGD algorithm, which updates the iterate as $w^{k+1}=w^{k}-\gamma\nabla f(w^{k};\eta^{k}),$ where $\{\eta^{k}\}$ are i.i.d., $w^{0}\in\mathbb{R}^{d}$ is a starting point, $\gamma$ is a step size, and $\Delta:=f(w^{0})-f^{*}.$ Arjevani et al., (2022); Carmon et al., (2020) showed that Vanilla SGD is optimal in terms of oracle complexity, which is given by $\Theta\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$ for finding an $\varepsilon$-stationary point.
Multiple workers and optimal time complexity. Consider now that we have $n$ workers computing stochastic gradients asynchronously and in parallel. In this setup, there are numerous ways to construct a distributed SGD method. The most well-known celebrated and recent approaches include Synchronized SGD (Minibatch SGD), Local SGD (Zinkevich et al.,, 2010; Stich,, 2019), Asynchronous SGD (Recht et al.,, 2011), Picky SGD (Cohen et al.,, 2021), Rennala SGD (Tyurin and Richtárik,, 2023), and Ringmaster ASGD (Maranjyan et al.,, 2025). The multi-worker setup is rich and versatile, offering numerous ways to design distributed SGD methods.
One may naturally ask which method offers the best theoretical performance. In distributed settings, the standard oracle complexity becomes less informative, as workers compute stochastic gradients in parallel with varying speeds. A more suitable comparison uses the $h_{i}$-fixed computation model (Mishchenko et al.,, 2022), where each worker $i$ needs at most $h_{i}$ seconds to compute a gradient. In this model, Mishchenko et al., (2022); Koloskova et al., (2022) showed that Asynchronous SGD outperforms Synchronized SGD. Its time complexity is further improved by Rennala SGD (Tyurin and Richtárik,, 2023) and Ringmaster ASGD [^1] Asynchronous SGD with a key modification; see Alg.7. (Maranjyan et al.,, 2025), both optimal under this and the more general universal computation model (Tyurin,, 2025) (see Section A). However, as we will discuss in more detail later, other factors come into play, such as communication complexity, support for AllReduce, peak bandwidth, and model update frequency.
@@ 1.2 Contributions
$\spadesuit$ New framework: Birch SGD (Section 2). We propose Birch SGD, a unifying framework that captures virtually all distributed SGD methods. The key idea is that SGD methods can be represented using weighted directed trees, which we refer to as computation trees (see Figure 1). We develop a new theoretical result, Theorem 2.4, that reduces the analysis of SGD methods to analyzing of the structure of these computation trees. The proofs become purely geometric and topological in nature, offering geometric intuition for the design of new methods.
$\clubsuit$ Eight new methods (Table 1 and Section 3). Using Birch SGD, we identify eight new methods in addition to those already known. For the first time, we prove that at least six of these newly discovered methods are computationally optimal, matching the lower bound (Tyurin and Richtárik,, 2023). We compare all methods across several dimensions, including computational and communication complexity, AllReduce compatibility, peak bandwidth, and model update frequency. Our improvements: i) our newly developed Async-Local SGD and Async-Batch SGD provably improve the communication complexity of Ringmaster ASGD while preserving asynchronicity; ii) we introduce Cycle SGD, which provably reduces peak bandwidth compared to all existing methods; iii) we propose a key modification to the family of local methods and design Local SGD and Dual-Process SGD, that, for the first time in the literature, achieve the optimal time complexities within this family; iv) for multi-cluster settings, we introduce Local-Async SGD and Nested Local-Async SGD, incorporating a carefully designed synchronization mechanism that guarantees optimality in computational time complexity; v) we develop a flexible meta-algorithm, Meta Local SGD, which supports arbitrary synchronization strategies, while incorporating a “Hard Sync” mechanism to guarantee convergence rates and to temper overly chaotic synchronization. As a byproduct, we prove that frequent model updates of fully asynchronous methods can lead to faster convergence and improve optimal Rennala SGD.
@@ 2 Birch SGD: A General View of SGD Methods
     Compute the new point [^2] The generated $\{\eta\}$ values are not necessarily i.i.d. For instance, we allow the algorithms to reuse previously computed stochastic gradients. This is important for Local SGD and other methods.: $w^{k+1}=w_{\textnormal{base}}-\gamma\nabla f(w_{\textnormal{grad}};\eta),\quad\eta\sim\mathcal{D}_{\xi}$
@@
Note that any instance of Algorithm 1, after any steps, can be represented by a weighted directed tree $G=(V,E),$ called a computation tree, where $V$ is the set of computed points and $E$ is the set of edges with weights given by the stochastic gradients used to compute the new points.
Our main idea now is to take any computation tree $G$ and analyze its structure to provide convergence guarantees for the corresponding SGD method. Intuitively, the structure of the tree, e.g., number of branches, length of branches, the tree distance between $w_{\textnormal{grad}}$ and $w_{\textnormal{base}}$ in Alg. 1 when we calculate a new point should be related to the convergence speed of the method.
@@ 2.1 Main theoretical result on convergence rates
As an example, consider Figure 2, where $\textnormal{dist}(w^{9},w^{4})=\max\{4,2\}=4,$ because the common ancestor is $w^{0},$ the number of edges from $w^{9}$ to $w^{0}$ is $4,$ and the number of edges from $w^{4}$ to $w^{0}$ is $2$ It is left to define the representation of a point $y\in V.$
@@
Condition 3: There exists a constant $R\in[0,\infty]$ such that $\textnormal{dist}(x^{k},z^{k})\leq R$ for all $k\geq 0.$ Then $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(x^{k})\|^{2}\right]\leq\varepsilon$ for all $K\geq\frac{4(R+1)L\Delta}{\varepsilon}+\frac{8\sigma^{2}L\Delta}{\varepsilon^{2}}$ with step size $\gamma=\min\{\frac{1}{2L},\frac{1}{2RL},\frac{\varepsilon}{4\sigma^{2}L}\}.$
Assumptions 1.1, 1.2, and 1.3 are well-known and standard in the analysis of stochastic optimization methods (Lan,, 2020; Arjevani et al.,, 2022). Let us explain the conditions of the theorem.
@@ 3 Existing and New Algorithms: Summary and Comparison
where $n$ workers collaboratively calculate the batch of size $B$ (see Alg. 4). This method produces a computation tree constructed as follows: $x^{1}=x^{0}-\gamma\nabla f(x^{0};\xi^{0}),\dots,x^{B}=x^{B-1}-\gamma\nabla f(x^{0};\xi^{B-1}),x^{B+1}=x^{B}-\gamma\nabla f(x^{B};\xi^{B}),\dots,x^{2B}=x^{2B-1}-\gamma\nabla f(x^{B};\xi^{2B-1}),\dots,$ where $B$ is a batch size (see Figure 14) and $\{\xi^{k}\}$ are i.i.d. from $\mathcal{D}_{\xi}.$ Notice that the computation tree is equivalent to (2) because $x^{B}=w^{1},x^{2B}=w^{2},$ etc. Here, all conditions of Theorem 2.4 are satisfied for the main branch $\{x^{k}\}$ with the auxiliary sequence $\{(z^{k},\xi^{k})\}$ such that $z^{0}=\dots=z^{B-1}=x^{0},$ $z^{B}=\dots=z^{2B-1}=x^{B},$ etc, and $\xi^{0}=\eta^{0,0},\dots,\xi^{B-1}=\eta^{0,B-1},\xi^{B}=\eta^{1,0},$ etc. However, unlike Vanilla SGD, $R=B-1$ because $\textnormal{dist}(x^{0},z^{0})=0,\textnormal{dist}(x^{1},z^{1})=1,\dots,\textnormal{dist}(x^{B-1},z^{B-1})=B-1,\textnormal{dist}(x^{B},z^{B})=0,$ etc. Thus, the iteration rate is $\mathcal{O}\left(\nicefrac{{BL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right).$
@@
[…]
@@
where $\delta^{k}$ is a delay such that $\delta^{k}\leq B-1,$ where $B\geq 1$ is a hyperparameter (see Alg. 7). We take $x^{k}=w^{k}$ for all $k\geq 0.$ Thus, the corresponding auxiliary sequence is defined by $z^{k}=x^{k-\delta^{k}}\equiv w^{k-\delta^{k}}$ and $\xi^{k}=\eta^{k-\delta^{k}}_{i}$ for all $k\geq 0.$ Constructing the computation tree (Figure 15), we can show that the conditions of Theorem 2.4 hold with $R=\max_{k\geq 0}\delta^{k}\leq B-1$ and the iteration rate is $\mathcal{O}\left(\nicefrac{{BL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right).$
@@
Previously, we presented Rennala SGD and Ringmaster ASGD that can be analyzed using Theorem 2.4. This raises the question: Which method is most effective, and how should one choose the appropriate one? In the following sections, we discuss different factors one should consider when selecting a method, and present new algorithms. The discussion here is summarized in Table 1. Before we begin, it is important to note that the iteration complexity in Theorem 2.4 does not reflect the true wall-clock performance. Instead, it serves as an intermediate result used to derive the time complexities presented below.
@@
1. Computational complexity. One way to compare the methods is to analyze their time complexity under the $h_{i}$-fixed computation model (see Sec. 1.1, A, and F). With a proper choice of the corresponding parameters, i.e., $B=\max\{1,\lceil\nicefrac{{\sigma^{2}}}{{\varepsilon}}\rceil\}$, both Rennala SGD and Ringmaster ASGD are optimal in terms of wall-clock time with the time complexity $\Theta(\min_{m\in[n]}[(\nicefrac{{1}}{{m}}\sum_{i=1}^{m}\nicefrac{{1}}{{h_{i}}})^{-1}\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{m\varepsilon^{2}}}\right)])$ provided that communication times are negligible. In the worst-case scenario, on the “very bad function” (Arjevani et al.,, 2022), all these methods perform equally well. Next, we discuss the strengths and weaknesses of the methods that are not captured by the $h_{i}$-fixed computation model.
@@
2. Number of model updates. At the same time, comparing (3) and (2) reveals that Rennala SGD computes $B$ stochastic gradients at the same point, while Ringmaster ASGD both computes and “explores” more by immediately updating the model as in (3). This feature makes Ringmaster ASGD more practically appealing. In fact, this intuition can be formalized using a simple two-dimensional strongly convex quadratic function $f:\mathbb{R}^{2}\to\mathbb{R}$ such that $f(x,y)=\mu x^{2}/2+Ly^{2}/2$ for all $x,y\in\mathbb{R}.$ For this function, we prove that Rennala SGD requires $\widetilde{\Theta}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\nicefrac{{\sigma^{2}}}{{\varepsilon n}}}\times h\times\nicefrac{{L}}{{\mu}}\right)$ seconds to achieve an $\varepsilon$–solution under the $h_{i}$-fixed computation model with $h_{i}=h$ for all $i\in[n]$, while Ringmaster ASGD needs $\widetilde{\Theta}\left(h\times\nicefrac{{L}}{{\mu}}\right),$ which is $\nicefrac{{\sigma^{2}}}{{\varepsilon n}}$ seconds less (see formal result in Section I). This is the first result showing that Ringmaster ASGD/Asynchronous SGD can be strongly better than Rennala SGD.

### v3

@@ 1 Introduction
Optimization is central to machine learning (ML), data science (DS), and federated learning (FL) (Konečný et al., 2016; Bottou et al., 2018; Kairouz et al., 2021). In these domains, stochastic optimization techniques such as stochastic gradient descent (SGD) (Robbins and Monro, 1951) and its adaptive variants (Adam, AdamW, etc) (Kingma and Ba, 2015; Loshchilov and Hutter, 2019) have become the standard approach for tackling large-scale problems (Schmidt et al., 2021). Due to the rising computational demands of modern functions, the theoretical foundation of distributed algorithms supporting a large number of workers (e.g., CPUs, GPUs, servers) is important (Mayer and Jacobsen, 2020; Kairouz et al., 2021; Douillard et al., 2023).
@@
In nonconvex settings, the goal is to find an $\varepsilon$–stationary point, meaning we want to find a random vector $\bar{x}$ such that $\mathbb{E}\left[\|\nabla f(\bar{x})\|^{2}\right]\leq\varepsilon$ (Nemirovskij and Yudin, 1983; Murty and Kabadi, 1985). The function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfies the following standard assumptions:
@@
$f$ is differentiable and LL–smooth: $\left\|\nabla f(x)-\nabla f(y)\right\|\leq L\left\|x-y\right\|$ $\forall x,y\in\mathbb{R}^{d}.$
@@
We focus on problems where workers are limited to computing stochastic gradients. Each worker has access to unbiased stochastic gradients, denoted by $\nabla f(x;\xi)$, whose variance is bounded by $\sigma^{2}$. In the context of ML, this implies that all workers can access the same data, which is practical when training large language and computer vision models. In such scenarios, privacy is not a critical concern, and devices can sample data from the Internet or shared datasets (Goodfellow et al., 2016).
@@ 1.1 Related work
One worker and optimal oracle complexity. With a single worker, the most standard optimization method is the Vanilla SGD algorithm, which updates the iterate as $w^{k+1}=w^{k}-\gamma\nabla f(w^{k};\eta^{k}),$ where $\{\eta^{k}\}$ are i.i.d., $w^{0}\in\mathbb{R}^{d}$ is a starting point, $\gamma$ is a step size, and $\Delta:=f(w^{0})-f^{*}.$ Arjevani et al. (2022); Carmon et al. (2020) showed that Vanilla SGD is optimal in terms of oracle complexity, which is given by $\Theta\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right)$ for finding an $\varepsilon$-stationary point.
Multiple workers and optimal time complexity. Consider now that we have $n$ workers computing stochastic gradients asynchronously and in parallel. In this setup, there are numerous ways to construct a distributed SGD method. The most well-known celebrated and recent approaches include Synchronized SGD (Minibatch SGD), Local SGD (Zinkevich et al., 2010; Stich, 2019), Asynchronous SGD (Recht et al., 2011), Picky SGD (Cohen et al., 2021), Rennala SGD (Tyurin and Richtárik, 2023), and Ringmaster ASGD (Maranjyan et al., 2025). The multi-worker setup is rich and versatile, offering numerous ways to design distributed SGD methods.
One may naturally ask which method offers the best theoretical performance. In distributed settings, the standard oracle complexity becomes less informative, as workers compute stochastic gradients in parallel with varying speeds. A more suitable comparison uses the $h_{i}$-fixed computation model (Mishchenko et al., 2022), where each worker $i$ needs at most $h_{i}$ seconds to compute a gradient. In this model, Mishchenko et al. (2022); Koloskova et al. (2022) showed that Asynchronous SGD outperforms Synchronized SGD. Its time complexity is further improved by Rennala SGD (Tyurin and Richtárik, 2023) and Ringmaster ASGD [^1] Asynchronous SGD with a key modification; see Alg.7. (Maranjyan et al., 2025), both optimal under this and the more general universal computation model (Tyurin, 2025) (see Section A). However, as we will discuss in more detail later, other factors come into play, such as communication complexity, support for AllReduce, peak bandwidth, and model update frequency.
@@ 1.2 Contributions
$\spadesuit$ New framework: Birch SGD (Section 2). We propose Birch SGD, a unifying framework that captures a wide range of distributed SGD methods. The key idea is that SGD methods can be represented using weighted directed trees, which we refer to as computation trees (see Figure 1). We develop a new theoretical result, Theorem 2.4, that reduces the analysis of SGD methods to analyzing of the structure of these computation trees. The proofs become purely geometric and topological in nature, offering geometric intuition for the design of new methods. Moreover, this geometric viewpoint leads to tighter time complexity guarantees even for Local SGD (FedAvg) approaches (McMahan et al., 2017), as we illustrate in Section H.
$\clubsuit$ Eight new methods (Table 1 and Section 3). Using Birch SGD, we identify eight new methods in addition to those already known. For the first time, we prove that at least six of these newly discovered methods are computationally optimal, matching the lower bound (Tyurin and Richtárik, 2023). We compare all methods across several dimensions, including computational and communication complexity, AllReduce compatibility, peak bandwidth, and model update frequency. Our improvements: i) our newly developed Async-Local SGD and Async-Batch SGD provably improve the communication complexity of Ringmaster ASGD while preserving asynchronicity; ii) we introduce Cycle SGD, which provably reduces peak bandwidth compared to all existing methods; iii) we propose a key modification to the family of local methods and design Local SGD and Dual-Process SGD that, for the first time in the literature, achieve the optimal time complexities within this family and improve upon the classical approach (see Section H); iv) for multi-cluster settings, we introduce Local-Async SGD and Nested Local-Async SGD, incorporating a carefully designed synchronization mechanism that guarantees optimality in computational time complexity; v) we develop a flexible meta-algorithm, Meta Local SGD, which supports arbitrary synchronization strategies, while incorporating a “Hard Sync” mechanism to guarantee convergence rates and to temper overly chaotic synchronization. As a byproduct, we prove that frequent model updates of fully asynchronous methods can lead to faster convergence and improve optimal Rennala SGD.
@@ 2 Birch SGD: A General View of SGD Methods
  Compute the new point22footnotemark: 2: $w^{k+1}=w_{\textnormal{base}}-\gamma\nabla f(w_{\textnormal{grad}};\eta),\quad\eta\sim\mathcal{D}_{\xi}$
@@
Note that any instance of Algorithm 1, after any steps, can be represented by a weighted directed tree $G=(V,E),$ called a computation tree, where $V$ is the set of computed points and $E$ is the set of edges with weights given by the stochastic gradients used to compute the new points. Our main idea now is to take any computation tree $G$ and analyze its structure to provide convergence guarantees for the corresponding SGD method. Intuitively, the structure of the tree, e.g., number of branches, length of branches, the tree distance between $w_{\textnormal{grad}}$ and $w_{\textnormal{base}}$ in Alg. 1 when we calculate a new point should be related to the convergence speed of the method.
Example. Consider Local SGD from Figure 4. There, we illustrate two global steps of the method with $2$ workers. In the first round, they compute $M_{1}=2$ and $M_{2}=2$ local steps (first figure in Figure 4). During these steps, worker $1$ first calculates $\nabla f(x^{0};\eta^{0,0}_{1}),$ finds $z^{0,1}_{1}=x^{0}-\gamma\nabla f(x^{0};\eta^{0,0}_{1}),$ then calculates $\nabla f(z^{0,1}_{1};\eta^{0,1}_{1})$ and $z^{0,2}_{1}=x^{0}-\gamma\nabla f(z^{0,1}_{1};\eta^{0,1}_{1}).$ Similar steps are performed by worker $2.$ Then, via a parameter-server or AllReduce, LocalSGD aggregates the stochastic gradients and performs the global step $x^{0}-\gamma(\nabla f(x^{0};\eta^{0,0}_{1})+\nabla f(z^{0,1}_{1};\eta^{0,1}_{1})+\nabla f(x^{0};\eta^{0,0}_{2})+\nabla f(z^{0,1}_{2};\eta^{0,1}_{2}))$ to obtain the new global point $x^{4},$ from which the second global steps will start. The step of finding $x^{4}$ is equivalent to the steps $x^{1}=x^{0}-\gamma\nabla f(x^{0};\eta^{0,0}_{1}),$ $x^{2}=x^{1}-\gamma\nabla f(z^{0,1}_{1};\eta^{0,1}_{1}),$ $x^{3}=x^{2}-\gamma\nabla f(z^{0,1}_{1};\eta^{0,1}_{1}),$ and $x^{4}=x^{3}-\gamma\nabla f(z^{0,1}_{2};\eta^{0,1}_{2}).$ This is how we construct the second figure in Figure 4, which is a geometric representation of the first global step. Then, the workers compute $M_{1}=1$ and $M_{2}=3$ local steps, accordingly, and synchronize again (third figure in Figure 4) to find $x^{8},$ from which the third global steps will start.
@@ 2.1 Main theoretical result on convergence rates
As an example, consider Figure 2, where $\textnormal{dist}(w^{9},w^{4})=\max\{4,2\}=4,$ because the common ancestor is $w^{0},$ the number of edges from $w^{9}$ to $w^{0}$ is $4,$ and the number of edges from $w^{4}$ to $w^{0}$ is $2.$ It is left to define the representation of a point $y\in V.$
@@
Condition 3: There exists a constant $R\in[0,\infty]$ such that $\textnormal{dist}(x^{k},z^{k})\leq R$ for all $k\geq 0.$ Then $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\|\nabla f(x^{k})\|^{2}\right]\leq\varepsilon$ for all $K\geq\frac{4(R+1)L\Delta}{\varepsilon}+\frac{8\sigma^{2}L\Delta}{\varepsilon^{2}}$ with step size $\gamma=\min\{\frac{1}{2L},\frac{1}{2RL},\frac{\varepsilon}{4\sigma^{2}L}\},$ where $\Delta=f(x^{0})-f^{*}.$
Assumptions 1.1, 1.2, and 1.3 are well-known and standard in the analysis of stochastic optimization methods (Lan, 2020; Arjevani et al., 2022). Let us explain the conditions of the theorem.
@@
Proof novelties. In Section D.1, we outline the key novelties, challenges, and the intuition guiding our choice of conditions. Although our proof in Section D.2 is compact—which we view as a strength rather than a limitation—it unifies a broad class of methods and provides new insights. Notably, right at the beginning, we introduce a distinct approach to handling the staleness term $\|x^{k}-z^{k}\|,$ which naturally arises from the update $x^{k+1}=x^{k}-\gamma\nabla f(z^{k};\xi^{k})$ in asynchronous and local methods. This treatment fundamentally differs from prior work, as it analyzes staleness through geometric graph reasoning. Moreover, using our framework, we later present our version of Local SGD, which yields tighter guarantees compared to the classical Local SGD (see Sections 3 and H), further validating both our framework and proof technique.
@@ 3 Existing and New Algorithms: Summary and Comparison
where $n$ workers collaboratively calculate the batch of size $B$ (see Alg. 4). This method produces a computation tree constructed as follows: $x^{1}=x^{0}-\gamma\nabla f(x^{0};\xi^{0}),\dots,x^{B}=x^{B-1}-\gamma\nabla f(x^{0};\xi^{B-1}),x^{B+1}=x^{B}-\gamma\nabla f(x^{B};\xi^{B}),\dots,x^{2B}=x^{2B-1}-\gamma\nabla f(x^{B};\xi^{2B-1}),\dots,$ where $B$ is a batch size (see Figure 16) and $\{\xi^{k}\}$ are i.i.d. from $\mathcal{D}_{\xi}.$ Notice that the computation tree is equivalent to (2) because $x^{B}=w^{1},x^{2B}=w^{2},$ etc. Here, all conditions of Theorem 2.4 are satisfied for the main branch $\{x^{k}\}$ with the auxiliary sequence $\{(z^{k},\xi^{k})\}$ such that $z^{0}=\dots=z^{B-1}=x^{0},$ $z^{B}=\dots=z^{2B-1}=x^{B},$ etc, and $\xi^{0}=\eta^{0,0},\dots,\xi^{B-1}=\eta^{0,B-1},\xi^{B}=\eta^{1,0},$ etc. However, unlike Vanilla SGD, R=B1R=B-1 because $\textnormal{dist}(x^{0},z^{0})=0,\textnormal{dist}(x^{1},z^{1})=1,\dots,\textnormal{dist}(x^{B-1},z^{B-1})=B-1,\textnormal{dist}(x^{B},z^{B})=0,$ etc. Thus, the iteration rate is $\mathcal{O}\left(\nicefrac{{BL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right).$
@@
Method (Sec. E) Optimal Computational Complexity (Sec. F) Communication Com-(g) plexity with Equal Times Optimal Total Complexity (Sec. I) AllReduce(a) (d)Update(d) Frequency (e)Peak(e) Bandwidth Rennala SGD (Alg. 4) (Tyurin and Richtárik, 2023) $\tau\frac{L\Delta}{\varepsilon}$ $\uparrow$ $n$ Ringmaster ASGD (Alg. 7) (Maranjyan et al., 2025) $\tau\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}}\vee\frac{L\Delta}{\varepsilon}\right)$ $\uparrow\uparrow\uparrow$ $n$ Local SGD (Alg. 5) (new)(f) $\tau\frac{L\Delta}{\varepsilon}$ $\uparrow\uparrow$ $n$ Cycle SGD (Alg. 8) (new) $\tau\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}}\vee\frac{L\Delta}{\varepsilon}\right)$ $\uparrow\uparrow$ $\frac{n^{2}\varepsilon}{\sigma^{2}}\vee 1$ Async-Local/Batch SGD (Alg. 9 and Sec. E.7) (new) $\tau\frac{L\Delta}{\varepsilon}$ $\uparrow\uparrow$ $n$ (Nested) Local-Async SGD (Alg. 12 and 14) (new) $(c)$ $(c)$ $\uparrow\uparrow$ $(c)$ Dual-Process SGD (Alg. 18) (new) $\tau\frac{L\Delta}{\varepsilon}$ $\uparrow\uparrow$ $n$ Meta Local SGD(b) (Alg. 16) (new) (a) Does a method support AllReduce? Asynchronous SGD-like methods do not support it due to their greedy update nature. Cycle SGD synchronizes only a subset of workers; thus, we cannot say definitively. We also cannot say definitively in the case of Local-Async SGD because the local asynchronous steps cannot be implemented with AllReduce, while the global steps can be. (b) This is an abstract method where all metrics (Computational Complexity, Communication Complexity, etc) depend on the chosen strategy. (c) Similar to AllReduce, here we also can not say definitely since Local-Async SGD and Nested Local-Async SGD are specially designed multi-cluster learning methods. (d) This is a slightly less formal metric that indicates how often an algorithm updates its iterate/model. Rennala SGD asks the workers to compute stochastic gradients at the same point; thus, it updates the iterates less frequently. In contrast, Ringmaster ASGD updates the iterates immediately. All other methods fall somewhere in between. See the discussion in Section 3. (e) In the Rennala SGD, Ringmaster ASGD, Local SGD, and Async-Local SGD methods, all workers can start communication simultaneously; thus, their peak bandwidth is $\mathcal{O}(n)$ when $n\leq\nicefrac{{\sigma^{2}}}{{\varepsilon}}$ In the Cycle SGD method, the workers communicate in a circular manner, so the peak bandwidth is $\mathcal{O}(\nicefrac{{n^{2}\varepsilon}}{{\sigma^{2}}})$ when $n\leq\nicefrac{{\sigma^{2}}}{{\varepsilon}}$, which is smaller. (f) While we recognize that Local SGD is well-known in the literature, what makes our version novel is the better time complexity compared to the classical version (Sec. H), the stopping criterion $\sum_{i=1}^{n}M_{i}=B$ in Alg. 5, and the analysis in Sec. E.3, F, and G, which leads to the optimal computational time complexities with a proper choice of $B.$ (g) We report the terms w.r.t. communication time $\tau$ under the $(h,\tau)$-fixed computation model from Section G.
@@
where $\delta^{k}$ is a delay such that $\delta^{k}\leq G-1,$ where $G\geq 1$ is a hyperparameter (see Alg. 7). We take $x^{k}=w^{k}$ for all $k\geq 0.$ Thus, the corresponding auxiliary sequence is defined by $z^{k}=x^{k-\delta^{k}}\equiv w^{k-\delta^{k}}$ and $\xi^{k}=\eta^{k-\delta^{k}}_{i}$ for all $k\geq 0.$ Constructing the computation tree (Figure 17), we can show that the conditions of Theorem 2.4 hold with $R=\max_{k\geq 0}\delta^{k}\leq G-1$ and the iteration rate is $\mathcal{O}\left(\nicefrac{{GL\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{\varepsilon^{2}}}\right).$
@@
Previously, we presented Rennala SGD and Ringmaster ASGD that can be analyzed using Theorem 2.4. This raises the question: Which method is most effective, and how should one choose the appropriate one? In the following sections, we discuss different factors one should consider when selecting a method, and present new algorithms. The discussion here is summarized in Table 1. Before we begin, it is important to note that the iteration complexity in Theorem 2.4 does not reflect the true wall-clock performance. It serves as an intermediate result used to derive the time complexities presented below.
@@
1. Computational complexity. One way to compare the methods is to analyze their time complexity under the $h_{i}$-fixed computation model (see Sec. 1.1, A, and F). With a proper choice of the corresponding parameters, i.e., $B=\max\{1,\lceil\nicefrac{{\sigma^{2}}}{{\varepsilon}}\rceil\}$, both Rennala SGD and Ringmaster ASGD are optimal in terms of wall-clock time with the time complexity $\Theta(\min_{m\in[n]}[(\nicefrac{{1}}{{m}}\sum_{i=1}^{m}\nicefrac{{1}}{{h_{i}}})^{-1}\left(\nicefrac{{L\Delta}}{{\varepsilon}}+\nicefrac{{\sigma^{2}L\Delta}}{{m\varepsilon^{2}}}\right)])$ provided that communication times are negligible. In the worst-case scenario, on the “very bad function” (Arjevani et al., 2022), all these methods perform equally well. Next, we discuss the strengths and weaknesses of the methods that are not captured by the $h_{i}$-fixed computation model.
@@
2. Number of model updates. At the same time, comparing (3) and (2) reveals that Rennala SGD computes $B$ stochastic gradients at the same point, while Ringmaster ASGD both computes and “explores” more by immediately updating the model as in (3). This feature makes Ringmaster ASGD more practically appealing. In fact, this intuition can be formalized using a simple two-dimensional strongly convex quadratic function $f:\mathbb{R}^{2}\to\mathbb{R}$ such that $f(x,y)=\mu x^{2}/2+Ly^{2}/2$ for all $x,y\in\mathbb{R}.$ For this function, we prove that Rennala SGD requires $\widetilde{\Theta}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\nicefrac{{\sigma^{2}}}{{\varepsilon n}}}\times h\times\nicefrac{{L}}{{\mu}}\right)$ seconds to achieve an $\varepsilon$–solution under the $h_{i}$-fixed computation model with $h_{i}=h$ for all $i\in[n]$, while Ringmaster ASGD needs $\widetilde{\Theta}\left(h\times\nicefrac{{L}}{{\mu}}\right),$ which is $\nicefrac{{\sigma^{2}}}{{\varepsilon n}}$ seconds less (see formal result in Section J). This is the first result showing that Ringmaster ASGD/Asynchronous SGD can be strongly better than Rennala SGD.